# Non-Resolution Reasoning (NRR): A Computational Framework for Contextual Identity and Ambiguity Preservation

### v6

@@ Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We illustrate NRR’s potential through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide functional verification in a synthetic two-turn disambiguation task through a Turn 1 Entropy measurement demonstrating ambiguity preservation directly: NRR-lite maintains high entropy ($H=0.63$, near-maximum $0.69$) at the ambiguous turn, while standard architectures collapse early ($H=0.10$), demonstrating that NRR preserves interpretive flexibility until context arrives. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
@@ 1.1 The Problem of Premature Semantic Collapse
Large language models (LLMs) have achieved remarkable success in natural language processing, demonstrating fluency in text generation, translation, and question-answering. However, these systems share a fundamental limitation: they over-commit to early resolution of ambiguity. This “premature semantic collapse bias” manifests as:
@@
The root cause lies in classical identity assumptions embedded in neural architectures. Standard embedding layers assign each token a fixed vector (or context-dependent but ultimately collapsed representation). Attention mechanisms use softmax normalization, forcing probabilistic collapse to sum-to-one distributions. Generation objectives optimize for single “correct” outputs rather than maintaining interpretation flexibility.
@@ 1.2 Contributions of This Work
    Functional Verification: We validate ambiguity preservation directly via Turn 1 Entropy measurement, showing NRR-lite maintains $H=0.63$ (near-maximum) while the baseline collapses to $H=0.10$, consistently across five random seeds.
@@ 5 Implementation Framework
This section presents architectural components for implementing NRR. While related to prior work on multi-prototype embeddings and non-softmax attention, the mechanisms in NRR are unified toward explicit ambiguity preservation rather than sense selection or sparsity.
@@ 5.2 Non-Collapsing Attention
Standard attention uses softmax normalization, forcing probabilistic collapse. NRR proposes non-collapsing functions (e.g., sigmoid) to allow multiple positions to receive strong attention simultaneously without competition.
@@ 4. Aggregate:
Stability could be maintained via gradient clipping and attention strength regularization.
@@ 5.5.2 Interference Operators
A typical pipeline: $\mathcal{S}^{\prime}=\pi(\kappa(\delta(\rho(\sigma(\mathcal{S}))),\mathcal{S}_{\text{new}}))$—stripping biases, positioning, dampening, integrating new information, and persisting.
@@ 6.3 Results
Model Turn 1 Entropy $H$ Gate Entropy Context Accuracy
Baseline $0.102\pm 0.089$ 100%
NRR-lite $\mathbf{0.629\pm 0.025}$ $0.693$ 100%
Maximum $H$ $0.693$ $0.693$
Figure 2: Turn 1 Entropy Verification Results. Left: Entropy at Turn 1 (no context). NRR-lite maintains high entropy ($H\approx 0.63$), near the theoretical maximum, while Baseline collapses ($H\approx 0.10$). Center: Distribution of $P(\text{FINANCIAL})$ at Turn 1. NRR-lite centers around 0.5 (ambiguous); Baseline clusters near 0 or 1 (committed). Right: Both models achieve 100% accuracy when Turn 2 context is provided, confirming that NRR-lite’s ambiguity preservation does not impair resolution capability.
@@
    Baseline Collapse ($H=0.10$): The baseline commits to one interpretation at Turn 1, even without disambiguating context. This premature collapse is the architectural bias NRR addresses.
@@
    NRR-lite Preservation ($H=0.63$): NRR-lite maintains near-maximum entropy at Turn 1, confirming that multiple interpretations coexist until context arrives.
@@
    Gate Behavior: NRR-lite’s context gate achieves $H=0.693$ (perfect balance) when no context is provided, shifting appropriately when Turn 2 arrives.
@@ 8 Conclusion
This paper introduced Non-Resolution Reasoning (NRR), a framework challenging classical identity assumptions in AI. In our controlled setting, we showed that preserving ambiguity can function as a capability rather than a cost. Our Turn 1 Entropy verification showed that NRR-lite maintains high interpretive entropy ($H=0.63$) when context is absent, while standard architectures collapse prematurely (H=0.10H=0.10)—yet both achieve perfect accuracy when context arrives. NRR challenges the assumption that meaning must collapse to be useful. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.

### v7

# NRR-Core: Non-Resolution Reasoning as a Computational Framework for Contextual Identity and Ambiguity Preservation

@@ Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We illustrate NRR’s potential through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide functional verification in a synthetic two-turn disambiguation task through a Turn 1 Entropy measurement demonstrating ambiguity preservation directly: NRR-lite maintains high entropy ($H=0.91$ bits, near-maximum $1.0$) at the ambiguous turn, while standard architectures collapse early ($H=0.15$ bits), demonstrating that NRR preserves interpretive flexibility until context arrives. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
@@ 1.1 The Problem of Premature Semantic Collapse
Large language models (LLMs) have achieved remarkable success in natural language processing [1, 3], demonstrating fluency in text generation, translation, and question-answering. However, these systems share a fundamental limitation: they over-commit to early resolution of ambiguity. This “premature semantic collapse bias” manifests as:
@@
The root cause lies in classical identity assumptions embedded in neural architectures. Standard embedding layers assign each token a fixed vector (or context-dependent but ultimately collapsed representation [2]). Attention mechanisms use softmax normalization [13], forcing probabilistic collapse to sum-to-one distributions. Generation objectives optimize for single “correct” outputs rather than maintaining interpretation flexibility. While recent decoding strategies introduce diversity at the output level [9, 10], the underlying representations still collapse internally.
@@
Several lines of prior work address related aspects of these limitations without targeting the core collapse problem. Efficient attention variants [8, 16] reduce computational cost but retain softmax normalization and its inherent collapse pressure. Mixture-of-experts architectures [7] maintain parallel computational pathways but route inputs to single experts, enforcing collapse at the routing level. Probabilistic programming languages [11] maintain distributions over hypotheses but typically aim for eventual convergence to point estimates. Neurosymbolic approaches [12] integrate logical and neural reasoning but inherit classical identity assumptions from their symbolic components. NRR differs from all of these by treating non-resolution—the sustained coexistence of conflicting interpretations—as a first-class computational primitive.
@@ 1.2 Contributions of This Work
    Functional Verification: We validate ambiguity preservation directly via Turn 1 Entropy measurement, showing NRR-lite maintains $H=0.91$ bits (near-maximum) while the baseline collapses to $H=0.15$ bits, consistently across five random seeds.
@@ 5 Implementation Framework
This section presents architectural components for implementing NRR. While related to prior work on multi-prototype embeddings [4, 5, 6] and non-softmax attention [14, 15], the mechanisms in NRR are unified toward explicit ambiguity preservation rather than sense selection or sparsity.
@@ 5.2 Non-Collapsing Attention
Standard attention uses softmax normalization, forcing probabilistic collapse. NRR proposes non-collapsing activation functions (e.g., sigmoid or other non-competitive functions) to allow multiple positions to receive strong attention simultaneously without competition.
@@ 4. Aggregate:
Stability could be maintained through techniques such as layer normalization [18] and residual scaling [17], along with gradient clipping and attention strength regularization.
@@ 5.5.2 Interference Operators
One possible pipeline: $\mathcal{S}^{\prime}=\pi(\kappa(\delta(\rho(\sigma(\mathcal{S}))),\mathcal{S}_{\text{new}}))$—stripping biases, positioning, dampening, integrating new information, and persisting.
Note on operator constraints: While NRR permits non-resolution as a valid state, not all transformations preserve this property. Operators must satisfy design principles (e.g., relative structure preservation) to avoid collapsing the very ambiguity NRR maintains. The formal specification of these constraints is addressed in subsequent work.
@@ 6.3 Results
Model Turn 1 Entropy $H$ (bits) Gate Entropy Context Accuracy
Baseline $0.15\pm 0.13$ 100%
NRR-lite $\mathbf{0.91\pm 0.04}$ $1.00$ 100%
Maximum $H$ $1.00$ $1.00$
Figure 2: Turn 1 Entropy Verification Results. Left: Entropy at Turn 1 (no context). NRR-lite maintains high entropy ($H\approx 0.91$ bits), near the theoretical maximum, while Baseline collapses ($H\approx 0.15$ bits). Center: Distribution of $P(\text{FINANCIAL})$ at Turn 1. NRR-lite centers around 0.5 (ambiguous); Baseline clusters near 0 or 1 (committed). Right: Both models achieve 100% accuracy when Turn 2 context is provided, confirming that NRR-lite’s ambiguity preservation does not impair resolution capability.
@@
    Baseline Collapse ($H=0.15$ bits): The baseline commits to one interpretation at Turn 1, even without disambiguating context. This premature collapse is the architectural bias NRR addresses.
@@
    NRR-lite Preservation ($H=0.91$ bits): NRR-lite maintains near-maximum entropy at Turn 1, confirming that multiple interpretations coexist until context arrives.
@@
    Gate Behavior: NRR-lite’s context gate achieves $H=1.0$ bit (perfect balance) when no context is provided, shifting appropriately when Turn 2 arrives.
@@ 8 Conclusion
This paper introduced Non-Resolution Reasoning (NRR), a framework challenging classical identity assumptions in AI. In our controlled setting, we showed that preserving ambiguity can function as a capability rather than a cost. Our Turn 1 Entropy verification showed that NRR-lite maintains high interpretive entropy ($H=0.91$ bits) when context is absent, while standard architectures collapse prematurely ($H=0.15$ bits)—yet both achieve perfect accuracy when context arrives. NRR challenges the assumption that meaning must collapse to be useful. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.